ul capacity of 4,200 sets per year. Annual cost dats at full capacity follow 104.000 107000 42.000 Product Cost factory office salaries
2. Compute the average product cost of one patio set (Round your answer to nearest whole dollar.) 3. Assume that production drops to only 1,000 sets annually. Would you expect the average product cost per set to increase, decrease, or remain unchanged? ゜Increase O Decrease
Answer #1

1)

Cost Behavior

Selling or Administrative

Product Cost

Variable

Fixed

Cost

Direct

Indirect

Direct labor

88,000

88,000

Advertising

104,000

104,000

Factory supervision

73,000

73,000

Property taxes, factory building

25,000

25,000

Sales commissions

63,000

63,000

Insurance, factory

7,000

7,000

Depreciation, administrative office equipment

2,000

2,000

Lease cost, factory equipment

13,000

13,000

Indirect materials, factory

19,000

19,000

Depreciation, factory building

108,000

108,000

Administrative office supplies (billing)

3,000

3,000

Administrative office salaries

107,000

107,000

Direct materials used (wood, bolts, etc.)

435,000

435,000

Utilities, factory

42,000

42,000

650,000

439,000

279,000

523,000

287,000

 

2) Solution: 192.86

Working:

Average product cost = (Direct + Indirect) / No of units = (523000 + 287,000) / 4200 = 192.86

3) Solution: Increase

Working: The average product cost per set will rise if the production increases because the fixed costs would be spread leading the average cost per unit to increase.